I need workings
aat Seller hree worKavc in a Competihe narkat Tha raw imaterial equived mak aach wot Costs 14 36 meat Selur hire tftk mast Seller sell hre?
Answer #1

-Here we need to find total cost which includes cost of raw material and labor cost

-Find MC for each labor hired

-Find MR for each labor hired

-The meat seller needs to hire till MR=MC and needs to stop thereafter

a)

No of workers

Meat Production/Day

Cost of raw material

Cost of labor at 80 per day

TC

MC

TR

MR

0

0

0

0

0

0

1

16

80

80

160

160

160

160

2

36

180

160

340

180

360

200

3

54

270

240

510

170

540

180

4

70

350

320

670

160

700

160

5

84

420

400

820

150

840

140

6

96

480

480

960

140

960

120

The meat seller needs to hire 4 number of workers

b.

No of workers

Meat Production/Day

Cost of raw material

Cost of labor at 70 per day

TC

MC

TR

MR

0

0

0

0

0

0

1

16

80

70

150

150

160

160

2

36

180

140

320

170

360

200

3

54

270

210

480

160

540

180

4

70

350

280

630

150

700

160

5

84

420

350

770

140

840

140

6

96

480

420

900

130

960

120

The meat seller needs to hire 5 number of workers

c.

No of workers

Meat Production/Day

MP

VP

MVP

0

0

0

0

1

16

16

128

128

2

36

20

288

160

3

54

18

432

144

4

70

16

560

128

5

84

14

672

112

6

96

12

768

96

Value product = Total product*price

4th labor Value product = 70*8=560

4th labor marginal value product = MP of 4th labor*price = 16*8=128

d.

No of workers

Meat Production/Day

Cost of raw material

Cost of labor at 80 per day

TC

MC

TR

MR

0

0

0

0

0

0

1

16

80

80

160

160

128

128

2

36

180

160

340

180

288

160

3

54

270

240

510

170

432

144

4

70

350

320

670

160

560

128

5

84

420

400

820

150

672

112

6

96

480

480

960

140

768

96

As MC is higher than MR, the meat seller needs to hire 0 labors, as it would leads to losses